OpenStudy (anonymous):

Would someone mind helping me with this quick precalc question? I have no idea how to solve this! Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164 Here are my choices: an = 896 - 53(n - 2) an = 896 - 53(n - 1) an = 896 + 53(n + 1) an = 896 - 53(n + 1)

OpenStudy (anonymous):

I've never worked with a problem like this before...

OpenStudy (loser66):

for option a) we test like: \(a_\color{red}{n}= 896-53 (\color{red}{n}-2)\) \(a_\color{red}{19} = 896 -53(\color{red}{19}-2)\) You calculate \( a_{19}=??\)

OpenStudy (anonymous):

I got -5 for a19

OpenStudy (anonymous):

Oh! So a must not be the answer because a19 is supposed to equal -58

OpenStudy (loser66):

Yes, it is -5 for \(a_{19}\) of that formula while our \(a_{19} = -58\)

OpenStudy (loser66):

ok, test option b show me your understanding.

OpenStudy (anonymous):

an = 896 - 53(n - 1) a19 = 896 - 53(19-1) a19=-58 B is my answer!

OpenStudy (loser66):

have to test \(a_{21}\) also, to make sure that you are not lucky when getting it.

OpenStudy (anonymous):

Yes, i tested a21 and got 896 - 53(21-1) and it was -164

OpenStudy (loser66):

Bingo!! congrats

OpenStudy (anonymous):

Thank you so much for helping me understand this! Really appreciate it!

OpenStudy (loser66):

np